A truck travelling at 22.5m/ decelerates at 2.27m/s².

How far does it travel during the third second when brakes are applied?


1
Expert's answer
2022-06-07T17:36:43-0400
"d(t)=v_0t+at^2\/2"


"d=d(3)-d(2)\\
=(v_0*3+a*3^2/2)-(v_0*2+a*2^2/2)\\
=v_0*1+5a*1^2/2\\
=22.5+5*(-2.27/2)=16.8\;\rm m"